Experimental values        

[S](mM) V (umol S / ml*sec)

2 1.19

4 13.9

What is Km of the enzyme?
Answer #1

The values ​​of 1 / [S] versus 1 / V are plotted to obtain the Lineweaver-Burk diagram:

Where:

1 / Vmax = b = 0.696

Vmax = 1 / 0.696 = 1.437

Y:

Km = m * Vmax = 3,072 * 1,437 = 4,414

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.